OpenStudy (anonymous):

HELP PLEASE!!! Sam Seller pays a local club $40 a month. Each week he has dinner for a flat fee of $20.00. Sam takes a cab, which charges him $1.75 for the first half-mile and $1.00 for each mile after that. The club is 5.5 miles from Sam's home, and he tips the driver $5.00 when he returns home. What does this meal cost Sam? (Don't forget the club membership for that week.) A.)$38.50 B.)$48.50 C.)$58.50 D.)$78.50

OpenStudy (anonymous):

1. im not sure i understand the question correctly but i think that u gotta add the costs for the meal and the taxi ride home because its because of the meal he has to take the taxi. i think. so therefore you would find how much it costs for the taxi which is: 1.75 + (5 x 1) = $5.75 then add the tip so it equals $10.75 add the meal cost so that's $30.75 altogether.

OpenStudy (anonymous):

75 + (5 x 1) = $5.75 then add the tip so it equals $10.75 add the meal cost so that's $30.75 altogether.

OpenStudy (anonymous):

that what i got

OpenStudy (anonymous):

So I Would Pick A?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thank You To All Of You!!!

OpenStudy (cj49):

1.75 + (5 x 1)=6.75 v.v

OpenStudy (cj49):

6.75+5=11.75+40=51.75+20=71.75+6.75=78.50

OpenStudy (cj49):

@amandanewsomm 30.75 is not the correct ans

OpenStudy (anonymous):

@cj49 why did you add 6.75 twice?

OpenStudy (cj49):

he tips the driver when he returns home

OpenStudy (anonymous):

sorry got a little distracted but @cj49 you right 78.50 is correct again sorry I have my little one with me right now so I got a little bit distracted

OpenStudy (anonymous):

Ah

OpenStudy (cj49):

in the last line of the question they have mentioned it that (Don't forget the club membership for that week) membership fees =40+flat fee for the mean 20=60 so the ans has to be more than $60